Can anyone please help me to simplify the two Z1 and Z2 expressions using boolean identities? then help me to draw the logisgm circuit.

(POS) Boolean algebra Z1 = (X1∙X2∙~X3∙X4) ∙ (X1∙~X2∙X3∙~X4) ∙ (~X1∙X2∙X3∙X4) ∙ (~X1∙X2∙3∙X4) ∙ (~X1∙~X2∙~X3∙X4) ∙ (~X1∙~X2∙~X3∙~X4)

Kmap simplify algebra: Z1 = (X2.~X3.X4)+(~X1.X2.X4)+(~X1.~X2.~X3)+(X1.~X2.X3.~X4)

(POS) Boolean algebra Z2 = (X1∙X2∙X3∙X4) ∙ (X1∙X2∙X3∙~X4) ∙ (X1∙~X2∙X3∙~X4) ∙ (X1∙~X2∙~X3∙~X4) ∙ (~X1∙X2∙X3∙~X4) ∙ (~X1∙X2∙~X3∙~X4) ∙ (~X1∙~X2∙X3∙~X4)

Kmap simplify algebra: Z1 = (X3.~X4)+(X1.X2.X3)+(X1.~X2.~X4)+(~X1.X2.~X4)

please help me to simplify these two;

Z1 = (X2.~X3.X4)+(~X1.X2.X4)+(~X1.~X2.~X3)+(X1.~X2.X3.~X4)

Z1 = (X3.~X4)+(X1.X2.X3)+(X1.~X2.~X4)+(~X1.X2.~X4)
Answer #1

DE Take any as 1 symbol example NX3 as to z = ( x + x2 + x3 +X4) (X* xo+ x2 + x4) (x + X2 + x 2 + x4) (x +X , fXg & Xy cxiX 2 + x3 + X4) (x1 + x2 + x3 +X4) cx + X2 + X₂ + X4) (x + X 2 + X3 & Xy) cxit x2 + x₂ + x4) ( x) +22² + X 3 + X, )))bemorgans law : (A+B+C+D) Alpic!! (ABCD]!= A+8!+c+D! +x, X2 X3 x yt x, X2 X₃ xy + x, X2 X 3 Xy) Idempotent Law A TAZA) 2Z2 = (x1+X at Xz+xy ) CX , tx g + X z w x y ) ( x tw X 2 + x zt w X4) (X, NX NXzt wxy) (NX + x + x + xy ) (NX,X tNX & X4) (NXJUU (2])=(((x + x2 + x3 + xy) (X1 + Xg+ *g+ xy!) (x, + xy + x3 + x4) ex, +x) +xz+xu (x,+ x2 + x3 + xyl) CX,+x2+xg+xy!) (X+ x, xox g xy + xi x;*xy + xy X2 xs xy) =(x) X, xxxy + xixx8xXy + x, X;\x3Xy ****y (w/x3+xx, 1 X2 =(xxx xx x + x,

Demorgans law : (A+B) = AB (AB) = AABI = ( x1 + x2 + x3) CX,+xy + x4) (x3 + x2 + x4) (x3 + x4) Replace an with N. =X, +